# Fixed Horizon Linear Quadratic Covariance Steering in Continuous Time with Hilbert-Schmidt Terminal Cost

### v1

@@ 1 Introduction
$\inf_{\bm{u}_{t}\in\mathcal{U}}\>\Phi\left(\bm{\Sigma}_{1},\bm{\Sigma}_{d}\right)+\mathbb{E}\left\{\!\int_{t_{0}}^{t_{1}}\!\!\left(\|\bm{u}_{t}\|_{2}^{2}+\bm{x}_{t}^{\top}\bm{Q}_{t}\bm{x}_{t}\right){\rm{d}}t\right\}\!$ (1a)
@@
where the time horizon is finite and fixed as $t_{0},t_{1}]$, the symbol $\mathbb{E}$ denotes the expectation w.r.t. the state vector $\bm{x}_{t}\in\mathbb{R}^{n}$, the $\mathcal{N}$ denotes multivariate normal probability density function (PDF), and $\sim$ is a shorthand for “follows the distribution”.
@@
with deterministic gain $\bm{K}_{t}:[t_{0},t_{1}]\mapsto\mathbb{R}^{m\times n}$, so as to steer the state covariance from $\bm{\Sigma}_{0}$ at $t=t_{0}$, to a neighborhood of $\bm{\Sigma}_{d}$ at $t=t_{1}$, that is optimal w.r.t. (1).
@@
The covariance $\bm{\Sigma}_{0}$ captures the initial condition uncertainty, and is usually available from an estimator. A terminal cost $\Phi\left(\bm{\Sigma}_{1},\bm{\Sigma}_{d}\right)$ allows the control designer to trade off the cost for terminal second-order statistics mismatch with the average cost-to-go $\mathbb{E}\left\{\!\int_{t_{0}}^{t_{1}}\!\!\left(\|\bm{u}_{t}\|_{2}^{2}+\bm{x}_{t}^{\top}\bm{Q}_{t}\bm{x}_{t}\right){\rm{d}}t\right\}$.
@@ 2 Conditions for Optimality
$H\left(\bm{\Sigma}_{t},\bm{P}_{t},\bm{K}_{t}\right)=\langle\bm{\Sigma}_{t},\bm{K}_{t}\bm{K}_{t}^{\top}+\bm{Q}_{t}\rangle$
@@
$+\langle\bm{\Sigma}_{t}+\bm{\mu}_{t}\bm{\mu}_{t}^{\top},\bm{K}_{t}\bm{K}_{t}^{\top}+\bm{Q}_{t}\rangle+2\bm{v}_{t}^{\top}\bm{K}_{t}\bm{\mu}_{t}+\bm{v}_{t}^{\top}\bm{v}_{t}$
@@
In light of Proposition (2), we are led to solve the coupled nonlinear system of equations:
@@ 3.1 Mappings $\bm{P}_{1}\mapsto\bm{P}_{0},\bm{H}_{0}\mapsto\bm{H}_{1}$
and its block-partitioned state transition matrix
@@
where $\bm{\Phi}_{ij}(s,t)\in\mathbb{R}^{n\times n}$ $\forall(i,j)\in\{1,2\}^{2}$. By definition, $\partial_{t}\bm{\Phi}(s,t)=\bm{M}(t)\bm{\Phi}(s,t)$, $\bm{\Phi}(s,s)=\bm{I}$.
@@ 3.2 Mapping $\bm{H}_{1}\mapsto\bm{P}_{1}$
For given $\bm{H}_{1}\in\mathbb{S}^{n},\bm{\Sigma}_{d}\succ\bm{0}$, it is unclear which of the non-unique symmetric solutions of (39) can be used for the mapping $\bm{H}_{1}\mapsto\bm{P}_{1}$, that can in turn guarantee convergence of the recursive algorithm.
@@ 4.1 Convergence
The proposed algorithm is a nonlinear recursion over the space of symmetric matrices. This recursion of the form $\bm{X}\mapsto\bm{F}(\bm{X})$, $\bm{X}\in\mathbb{S}^{n}$, where the matricial map $\bm{F}:\mathbb{S}^{n}\to\mathbb{S}^{n}$. Our main result is Theorem 4 that builds on Lemma 2 in Appendix .1. Specifically, Theorem 4 establishes the existence-uniqueness of the fixed point $\bm{X}^{\star}\in\mathbb{S}^{n}$ of the proposed recursion, and its convergence.
@@
Now we argue that $\bm{F}$ has at least one fixed point. To do so, we note that the set $\mathbb{S}^{n}$ is closed and convex (w.r.t. standard Euclidean norm topology). The composite map $\bm{F}:\mathbb{S}^{n}\to\mathbb{S}^{n}$ is continuous and bounded. This can be seen from the fact that by Lemma 2, $\bm{F}_{4}$ is bounded, and by similar argument, $\bm{F}_{2}$ is bounded as well. The map $\bm{F}_{1}$, being affine, is non-expansive. By Lemma 3, $\bm{F}_{3}$ is non-expansive as well. Finally, the $\bm{F}_{i}$ are continuous for all $i=1,\ldots,4$. So for any $\bm{X}\in\mathbb{S}^{n}$ with finite norm, the half-vectorization map $\bm{f}:={\mathrm{vech}}(\bm{F}(\bm{X})):\mathbb{R}^{n(n+1)/2}\to\mathbb{R}^{n(n+1)/2}$ is uniformly bounded by some finite $M>0$, i.e.,
@@ 5.1 Noisy double integrator
Figure 6: The relative orbital dynamics for the noisy Clohessy-Wiltshire model in Sec. 5.2.

### v2

@@ 1 Introduction
$\inf_{\bm{u}_{t}\in\mathcal{U}}\>\phi\left(\bm{\Sigma}_{1},\bm{\Sigma}_{d}\right)+\!\int_{t_{0}}^{t_{1}}\!\!\mathbb{E}\left(\|\bm{u}_{t}\|_{2}^{2}+\bm{x}_{t}^{\top}\bm{Q}_{t}\bm{x}_{t}\right){\rm{d}}t$ (1a)
@@
where the time horizon is finite and fixed as $t_{0},t_{1}]$, the symbol $\mathbb{E}$ denotes the expectation w.r.t. the state vector $\bm{x}_{t}\in\mathbb{R}^{n}$, the $\mathcal{N}$ denotes multivariate normal probability density function (PDF), and $\sim$ is a shorthand for “follows the distribution”. In (1c), the $\rho_{0}$ is the PDF for the initial state $\bm{x}_{0}$, and the $\rho_{d}$ is the desired PDF for the terminal state at $t=t_{1}$. Depending on the choice of control policy, the statistics of the controlled terminal state $\bm{x}_{1}:=\bm{x}(t=t_{1})$ may not match with the desired terminal statistics $\rho_{d}$.
@@
with deterministic time-varying gain $\bm{K}_{t}:[t_{0},t_{1}]\mapsto\mathbb{R}^{m\times n}$, so as to steer the state covariance from $\bm{\Sigma}_{0}$ at $t=t_{0}$, to a neighborhood of $\bm{\Sigma}_{d}$ at $t=t_{1}$, that is optimal w.r.t. (1).
@@
The covariance $\bm{\Sigma}_{0}$ captures the initial condition uncertainty, and is usually available from an estimator. A terminal cost $\phi\left(\bm{\Sigma}_{1},\bm{\Sigma}_{d}\right)$ allows the control designer to trade off the cost for terminal second-order statistics mismatch with the average cost-to-go $\int_{t_{0}}^{t_{1}}\mathbb{E}\left(\|\bm{u}_{t}\|_{2}^{2}+\bm{x}_{t}^{\top}\bm{Q}_{t}\bm{x}_{t}\right){\rm{d}}t$.
@@ 2 Conditions for Optimality
$H\left(\bm{\Sigma}_{t},\bm{P}_{t},\bm{K}_{t}\right)=\langle\bm{\Sigma}_{t},\bm{K}_{t}^{\top}\bm{K}_{t}+\bm{Q}_{t}\rangle$
@@
$+\langle\bm{\Sigma}_{t}+\bm{\mu}_{t}\bm{\mu}_{t}^{\top},\bm{K}_{t}^{\top}\bm{K}_{t}+\bm{Q}_{t}\rangle+2\bm{v}_{t}^{\top}\bm{K}_{t}\bm{\mu}_{t}+\bm{v}_{t}^{\top}\bm{v}_{t}$
@@
In light of Proposition 2, we are led to solve the coupled nonlinear system of equations:
@@ 3.1 Mappings $\bm{P}_{1}\mapsto\bm{P}_{0},\bm{H}_{0}\mapsto\bm{H}_{1}$
and its state transition matrix in block-partitioned form:
@@
where $\bm{\Phi}_{ij}(s,t)\in\mathbb{R}^{n\times n}$ $\forall(i,j)\in\{1,2\}^{2}$. By definition,
$\partial_{t}\bm{\Phi}(s,t)=\bm{M}(t)\bm{\Phi}(s,t),\quad\bm{\Phi}(s,s)=\bm{I}.$ (22)
@@ 3.2 Mapping $\bm{H}_{1}\mapsto\bm{P}_{1}$
For given $\bm{H}_{1}\in\mathbb{S}^{n},\bm{\Sigma}_{d}\succ\bm{0}$, it is unclear which of the non-unique symmetric solutions of (40) can be used for the mapping $\bm{H}_{1}\mapsto\bm{P}_{1}$, that can in turn guarantee convergence of the recursive algorithm. This analysis suggests that instead of removing the nonlinearity from (38)-(39), we should try to leverage the same.
@@ 4.1 Convergence
The proposed algorithm is a nonlinear recursion over the space of symmetric matrices. This recursion of the form $\bm{X}\mapsto\bm{F}(\bm{X})$, $\bm{X}\in\mathbb{S}^{n}$, where the matricial map $\bm{F}:\mathbb{S}^{n}\to\mathbb{S}^{n}$. Our main result is Theorem 4 that builds on auxiliary lemmas 2-4 in Appendix .1. Specifically, Theorem 4 establishes the existence-uniqueness of the fixed point $\bm{X}^{\star}\in\mathbb{S}^{n}$ of the proposed recursion, and its convergence.
@@
Now we argue that $\bm{F}$ has at least one fixed point. To do so, we note that the set $\mathbb{S}^{n}$ is closed and convex (w.r.t. standard Euclidean norm topology). The composite map $\bm{F}:\mathbb{S}^{n}\to\mathbb{S}^{n}$ is continuous and bounded. This can be seen from the fact that by Lemma 2 in Appendix .1, the map $\bm{F}_{4}$ is bounded, and by similar argument, $\bm{F}_{2}$ is bounded as well. The map $\bm{F}_{1}$, being affine, is non-expansive. By Lemma 3 in Appendix .1, the map $\bm{F}_{3}$ is non-expansive as well. Finally, the $\bm{F}_{i}$ are continuous for all $i=1,\ldots,4$. So for any $\bm{X}\in\mathbb{S}^{n}$ with finite norm, the half-vectorization map $\bm{f}:={\mathrm{vech}}(\bm{F}(\bm{X})):\mathbb{R}^{n(n+1)/2}\to\mathbb{R}^{n(n+1)/2}$ is uniformly bounded by some finite $M>0$, i.e.,
@@ 5.1 Noisy double integrator
Figure 6: The relative orbital dynamics for the noisy Clohessy–Wiltshire model in Sec. 5.2.